OpenStudy (anonymous):

Find all y such that the distance between the points (-6, -9) and (-4, y) is 28.

OpenStudy (anonymous):

Label your points (x1, y1) and (x2, y2). The distance between two points is \[d = \sqrt{\left( x_2-x_1 \right)^2 + (y_2-y_1)^2}\]

OpenStudy (jdoe0001):

\(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ % (a,b) &({\color{red}{ -6}}\quad ,&{\color{blue}{ -9}})\quad % (c,d) &({\color{red}{ -4}}\quad ,&{\color{blue}{ y_2}}) \end{array}\qquad % distance value d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2} \\ \quad \\ 28 = \sqrt{({\color{red}{ -4}}-{\color{red}{(-6) }})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ (-9)}})^2}\impliedby solve\ for\ {\color{blue}{ y_2}}\)

OpenStudy (anonymous):

@jdoe0001 i got -57